OpenStudy (anonymous):

Use identities to find the exact value of the four remaining trig functions: Sin=1/2, cos=square root 3/2 I have that the tan is square root 3/3 Can someone help me find the other four functions?

OpenStudy (dumbcow):

csc = 1/sin sec = 1/cos cot = 1/tan